OpenStudy (anonymous):

Help Needed! To make orange paint, you mix 7 litres of yellow 2ith 3 litres of red paint. How many litres of each color do you need to make 120 litres of orange paint?

OpenStudy (anonymous):

7 + 3 = 10 120 / 10 = 12 12*7 = 84 yellow 12*3 = 36 red

OpenStudy (anonymous):

Thanks Dude! :)